IDX Technologies is a privately held developer of advanced security systems based in Chicago. As part of your business development​ strategy, in late 2013 you initiate discussions with​ IDX's founder about the possibility of acquiring the business at the end of 2013. Estimate the value of IDX per share using a discounted FCF approach and the following​ data: bullet ​Debt: $ 21 million bullet Excess​ cash: $ 102 million bullet Shares​ outstanding: 50 million bullet Expected FCF in​ 2014: $ 44 million bullet Expected FCF in​ 2015: $ 50 million bullet Future FCF growth rate beyond​ 2015: 5 % bullet ​Weighted-average cost of​ capital: 9.4 %

The enterprise value in 2013 is ​$ ? million. ​ (Round to the nearest​ integer.) The equity value is ​$ ? million. ​ (Round to the nearest​ integer.) The value of IDX per share is ​$ ?. ​ (Round to the nearest​ cent.)
Answer #1

Value of business= FCF1/(1+r)^1 + …………FCFn/(1+r)^n+ (FCFn+1/ (k-g))/(1+r)^n

= 44/1.094^1+50/1.094^2+ (50*105%/(9.4%-5%))/1.094^2

= 1078.94 million

Equity = Value of business- Debt-Excess cash

= 1078.94-21-102

= 955.94 million

Value of Share = 955.94 m / 50 m

= $19.12